# T2VUnlearning: A Concept Erasing Method for Text-to-Video Diffusion Models

### v2

@@ Abstract
Recent advances in text-to-video (T2V) diffusion models have significantly enhanced the quality of generated videos. However, their ability to produce explicit or harmful content raises concerns about misuse and potential rights violations. Inspired by the success of unlearning techniques in erasing undesirable concepts from text-to-image (T2I) models, we extend unlearning to T2V models and propose a robust and precise unlearning method. Specifically, we adopt negatively-guided velocity prediction fine-tuning and enhance it with prompt augmentation to ensure robustness against LLM-refined prompts. To achieve precise unlearning, we incorporate a localization and a preservation regularization to preserve the model’s ability to generate non-target concepts. Extensive experiments demonstrate that our method effectively erases a specific concept while preserving the model’s generation capability for all other concepts, outperforming existing methods. We provide the unlearned models in https://github.com/VDIGPKU/T2VUnlearning.git.
@@ 1 Introduction
To address this threat, one might consider retraining T2V models with a filtered dataset. However, this can be extremely expensive and thus infeasible. As an alternative, current approaches rely on prompt manipulation techniques, such as negative prompting, which guides model generation away from undesirable content using a negative prompt, or SAFREE [32], which projects toxic tokens in prompts into a harmless subspace. While these methods may work for simple prompts, SOTA T2V models default to using LLM-refined prompts rich in details to guide generation, which significantly reduces the effectiveness of prompt manipulation. This leads us to consider unlearning [2], which offers a way for trained models to selectively forget knowledge or data without complete retraining. Unlearning has been extensively studied in text-to-image (T2I) models, typically via fine-tuning with pseudo data generated by the original model to reduce the probability of generating images containing the target concept [12]. However, such unlearning methods are not robust in the T2V setting. Experiments show that when a concept (e.g., “beagle dog”) is unlearned, the model can still generate accurate videos of the concept if the prompt includes additional contextual information(e.g., “a beagle dog running on a grassy lawn”).
Figure 1: We propose T2VUnlearning, a robust and precise unlearning method for T2V models. T2VUnlearning effectively prevents the generation of videos containing target concepts while preserving the model’s knowledge of non-target concepts. We manually mask the images (*) for publication.
@@
To alleviate issues above, we propose T2VUnlearning, a robust and precise unlearning method for T2V models. Specifically, T2VUnlearning minimizes the probability of the model generating videos about a target concept by negatively-guided velocity prediction. Similar to ESD [12], we leverage the original model’s knowledge to generate pseudo-training data. However, instead of directly using the target concept as the generation prompt, we utilize an uncensored LLaMA3 [15] model to combine the concept with rich context, effectively preventing the generation of the target concept even when prompted with additional context or details, thereby significantly enhancing unlearning robustness.
@@
However, introducing context into the pseudo data can inadvertently cause the model to forget related knowledge (e.g., unlearning “beagle dog” with videos of “a beagle dog running on grassy lawn” might also cause the model to forget “lawn”). Additionally, we observe that unlearning a single concept can cause the model to forget a variety of semantically related non-target concepts (e.g., removing “beagle dog” degrades the model’s ability to generate other animal classes). This phenomenon is similar to Catastrophic Forgetting (CF) [11], where a model forgets previously learned information when learning new knowledge. To achieve precise unlearning, we employ two regularizations: localization and preservation. The former is inspired by Receler [19] that uses cross-attention masks to localize and constrain unlearning to the target concept. Although SOTA T2V models do not employ cross-attention, their cross-modal full-attention layers exhibit similar text-visual interactions [3], enabling us to derive analogous attention masks for localized unlearning. The latter is inspired by DreamBooth [25] that mitigates CF by introducing a prior-preservation term. We adopt a similar approach: by selecting a semantically related concept and encouraging the model to retain its ability to generate it, we can significantly enhance the preservation of other non-target concepts.
@@
  • To the best of our knowledge, we are the first to propose a fine-tuning-based unlearning method for T2V models.
  • Through prompt augmentation, localization, and preservation regularization, our method effectively erases target concepts while minimizing the impact on non-target concepts.
  • Extensive experiments show that our method outperforms current prompt manipulation techniques across diverse prompt types and video generation models.
@@ 2 Related Work
While U-Net-based approaches achieved early success, they struggle with modeling long-range temporal dependencies and suffer from computational inefficiencies. To address these limitations, recent T2V models have shifted toward Diffusion Transformer (DiT) architectures and removed cross-attention. CogVideoX [31] adopts a Multimodal DiT (MMDiT)  [10] that concatenates text and visual tokens as input to a cross-modal 3D full-attention layer, improving both efficiency and performance. HunyuanVideo [22] further explores this direction by introducing dual-stream and single-stream DiT variants: the former processes text and visual tokens separately, while the latter feeds their concatenation into a full-attention layer.
@@
The second category is unlearning, which alters the model weights by zero-shot model editing or fine-tuning methods. Zero-shot approaches directly manipulate the model weights—typically within the cross-attention layers—to suppress undesired concepts, as in UCE [13] and RECE [14]. While efficient and easy to apply, these methods rely heavily on cross-attention mechanism, making them incompatible with recent T2V architectures. In contrast, fine-tuning-based methods offer a more generalizable solution. ESD [12] erases concepts by learning to predict noise guided by negative prompts. AdvUnlearn [33] and Receler [19] build upon ESD with adversarial strategies to improve erasure robustness. Despite the progress in the T2I domain, experimental results show that existing methods fail to achieve robust and precise concept erasure in the T2V setting.
@@ 3.1 Background
Diffusion models [17] learn to generate data by reversing a process that gradually adds Gaussian noise to input. Given an input data point $\mathbf{x}$ and condition $y$, a diffusion model $\epsilon_{\theta}$ is trained to predict the noise $\epsilon$ from a noisy sample $\mathbf{x}_{t}$ at a given time step $t$. The training objective is to minimize the L2 distance between the model prediction and the actual noise:
@@
In T2V models, this process operates on latent representations of videos. We focus on two commonly used open-source models: CogVideoX [31] and HunyuanVideo [22].
CogVideoX employs a v-prediction parameterization [26], where the model is trained to predict a velocity vector instead of the added noise. Specifically, given a video latent representation $\mathbf{x}$ and condition $y$, the model adds Gaussian noise $\epsilon$ by $t$ steps and the noisy sample can be expressed as $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon$, where $\alpha_{t}$ and $\sigma_{t}$ are scheduling coefficients. The model $\mathbf{v}_{\theta}$ is encouraged to predict the ground truth velocity vector $\mathbf{v}=\alpha_{t}\mathbf{x}-\sigma_{t}\epsilon$ by the following loss function:
@@
HunyuanVideo is based on the flow-matching framework [23], which trains the model to learn a vector field that maps a complex data distribution to a simple Gaussian distribution. Specifically, given a latent video representation $\mathbf{x}$ and condition $y$, the model first initializes Gaussian noise $\mathbf{x}_{0}$ and constructs a training sample $\mathbf{x}_{t}$ at step $t$ by linear interpolation between $\mathbf{x}_{0}$ and $\mathbf{x}$. The model $\mathbf{v}_{\theta}$ is trained to predict the velocity vector $\mathbf{u}_{t}=d\mathbf{x}_{t}/dt$ that guides the noise back to the data by the following loss function:
$\mathcal{L}_{\textrm{hunyuan}}=\mathbb{E}_{\mathbf{x},y,\mathbf{x}_{0},t}\left\|\mathbf{u}_{t}-\mathbf{v}_{\theta}\left(\mathbf{x}_{t},y,t\right)\right\|_{2}^{2}.$ (3)
@@
$P_{\theta^{{}^{\prime}}}(\mathbf{x})\propto\frac{P_{\theta}(\mathbf{x})}{P_{\theta}(c|\mathbf{x})},$ (4)
where the $\theta^{{}^{\prime}}$ indicate new model weights after unlearning. With diffusion models that is trained with $\epsilon$-prediction, this goal can be reparametrized to predicting negatively guided noise $\epsilon_{neg}$:
$\epsilon_{neg}$ $=\epsilon_{\theta}(\mathbf{x}_{t},t)-\eta(\epsilon_{\theta}(\mathbf{x}_{t},c,t)-\epsilon_{\theta}(\mathbf{x}_{t},t)),$ (5)
$\mathcal{L}_{\textrm{unlearn}}$ $=\mathbb{E}_{\mathbf{x},c,\epsilon,t}\left\|\epsilon_{neg}-\epsilon_{\theta^{{}^{\prime}}}\left(\mathbf{x}_{t},c,t\right)\right\|_{2}^{2},$
where $\eta$ is a tunable parameter indicating unlearning strength. Note $\epsilon_{neg}$ resembles negative guidance applied in classifier-free guidance [16], which means the model’s prediction is negatively guided to deviate from its original prediction conditioned on the concept $c$. Thus, this process is referred to as negatively-guided noise prediction.
@@ 3.2 Overview of Pipeline
We propose T2VUnlearning, an adapter-based fine-tuning method that achieves robust and precise unlearning through three strategies: negatively-guided velocity prediction with prompt augmentation, mask-based localization regularization, and concept preservation regularization. An overview of the pipeline is illustrated in Fig. 2.
Figure 2: Overview of T2VUnlearning. T2VUnlearning performs unlearning by inserting adapters after each cross-modal full-attention layer. It generates pseudo data through prompt augmentation, and fine-tunes the model using an unlearning loss $\mathcal{L}_{unlearn}$ to erase the target concept, a mask-based localization regularization $\mathcal{L}_{loc}$, and a preservation regularization $\mathcal{L}_{pre}$ to retain non-target concepts.
3.3 Negatively-guided Prediction with Prompt Augmentation
Given any T2V model $\mathbf{v}_{\theta}$ with weights $\theta$, our goal is to obtain an unlearned model weight $\theta^{{}^{\prime}}$ that reduces the probability of generating video $\mathbf{x}$ described by target concept $c$. Based on Eq. 15, we can reparameterize this goal into predicting the negatively-guided velocity $\mathbf{v}_{neg}$:
@@ 3.3 Negatively-guided Prediction with Prompt Augmentation
$\mathcal{L}_{unlearn}$ $=\mathbb{E}_{\mathbf{x},c,\epsilon,t}\left\|\mathbf{v}_{neg}-\mathbf{v}_{\theta^{{}^{\prime}}}\left(\mathbf{x}_{t},c,t\right)\right\|_{2}^{2}.$
A detailed derivation of this objective is provided in Appendix 2. Note that while the velocity vector is defined differently in CogVideoX and HunyuanVideo, the formulation of the unlearning loss remains consistent across both models. Therefore, we employ the same notation $\mathbf{v}$ for simplicity.
Following ESD [12], instead of manually collecting real training data, we can utilize the knowledge of the original T2V model to generate pseudo data. Specifically, to obtain the $\mathbf{x}_{t}$ at step $t$, it is generated by denoising a sampled Gaussian noise to timestep $t$ instead of adding noise to real data. Previous works typically generate $\mathbf{x}_{t}$ using the target concept as the prompt. However, training only on simple instances of concept $c$ is insufficient to reduce the generation probability for prompts with additional context or detail. This is because these more complex examples, while still can be described or classified as $c$, come from a different distribution that the model hasn’t seen during fine-tuning. Therefore, we propose a simple yet effective solution: augmenting the prompts used to generate $\mathbf{x}_{t}$, that is, following the prompt-rewriting strategies employed during the training of T2V models. To reduce computational overhead, we utilize an uncensored variant of LLaMA3***https://huggingface.co/Orenguteng/Llama-3-8B-Lexi-Uncensored and apply in-context prompting to generate augmented prompts that align with the format requirements of T2V models. Fig. 3 is an illustration of the prompt augmentation process and the generation of pseudo data $\mathbf{x_{t}}$ at timestep $t$. Experimental results (see Section 4.4) demonstrate that this approach achieves robust unlearning when tested against a wide range of paraphrased prompts, including those written by humans.
Figure 3: Illustration of prompt augmentation for HunyuanVideo.
@@ 3.4 Mask-based Localization
Incorporating contextual information into the pseudo data $\mathbf{x}$ can inadvertently cause the model to forget knowledge of those additional contexts and details (e.g., a model trained to unlearn “beagle dog” using videos of “a beagle dog running on a grassy lawn” may also lose knowledge of “lawn”). Prior work in T2I, such as  Huang et al. [19], addresses this issue by computing a concept-related cross-attention mask and restricting model updates to the masked region. However, this strategy is difficult to adapt directly to T2V models, as they typically do not employ text-vision cross-attention but cross-modal full attention.
Despite the absence of explicit cross-attention in T2V models, interactions between text and visual tokens still exist. As demonstrated by DiTCtrl [3], the query-key (QK) maps in full-attention layers can be divided into distinct regions, among which the text-to-video region exhibits interactions analogous to cross-attention. Consequently, extracting a concept-related attention mask based on the text-to-video region of the QK map is feasible. An example of obtaining such attention masks in HunyuanVideo is illustrated in Fig. 4.
Figure 4: We compute the concept-related attention map by averaging the T2V QK attention scores corresponding to the token indices of the target concept. The attention map is then thresholded to obtain the final concept mask.

### v3

@@ Abstract
Recent advances in text-to-video (T2V) diffusion models have significantly enhanced the quality of generated videos. However, their capability to produce explicit or harmful content introduces new challenges related to misuse and potential rights violations. To address this newly emerging threat, we propose unlearning-based concept erasing as a solution. First, we adopt negatively-guided velocity prediction fine-tuning and enhance it with prompt augmentation to ensure robustness against prompts refined by large language models (LLMs). Second, to achieve precise unlearning, we incorporate mask-based localization regularization and concept preservation regularization to preserve the model’s ability to generate non-target concepts. Extensive experiments demonstrate that our method effectively erases a specific concept while preserving the model’s generation capability for all other concepts, outperforming existing methods. We provide the unlearned models in https://github.com/VDIGPKU/T2VUnlearning.git.
@@ 1 Introduction
To address this new threat, one might consider retraining T2V models with a filtered dataset. However, this can be extremely expensive and thus infeasible. As an alternative, current approaches rely on prompt manipulation techniques, such as SAFREE [32] which projects toxic tokens in prompts into a harmless subspace. However, SOTA T2V models use LLM-refined prompts rich in details to guide generation by default, which significantly reduces the effectiveness of prompt manipulation.
Figure 1: Illustration of concept erasing results for HunyuanVideo. T2VUnlearning successfully prevent the generation of videos containing the specified target concepts while preserving the knowledge of non-target concepts. Explicit videos (*) have been manually masked for publication.
@@
To alleviate the above issues, we propose T2VUnlearning, a robust and precise concept erasing method for T2V models, which utilizes unlearning [2] to selectively forget knowledge of undesirable concepts. Firstly, we revise the negatively-guided prediction of Text-to-Image (T2I) unlearning methods [12] to accommodate the different generative paradigms of T2V models. Specifically, we fine-tune T2V models with negatively-guided velocity prediction to minimize the probability of generating videos of target concepts, and use prompt augmentation to enhance robustness against LLM refined prompts. Unlike [12] that directly uses the target concept as the generation prompt to generate pseudo-training data, we exploit an uncensored LLaMA3 model [15] to augment the generation prompt with additional context or details, effectively prevent T2V models from generating undesirable contents even when prompted with LLM-refined prompts.
Secondly, we observe that introducing context into the pseudo data can inadvertently cause the model to forget context-related knowledge (e.g., unlearning “dog” with videos of “a dog running on grassy lawn” might also cause the model to forget “lawn”). Hence, we introduce mask-based localization regularization to address this issue. Specifically, we derive attention masks from the text-visual interaction areas of full-attention layers in T2V models [3], and use these masks to localize and constrain unlearning to only affect the target concept like [18].
@@
Furthermore, we observe that unlearning a single concept can cause the model to forget a variety of semantically related non-target concepts (e.g., removing “dog” degrades the model’s ability to generate other animal classes). This phenomenon is similar to Catastrophic Forgetting (CF) [11], where a model forgets previously learned information when learning new knowledge. To prevent the model from forgetting non-target concepts, we propose concept preservation regularization which is inspired by DreamBooth [25]. Specifically, by selecting a semantically related preservation concept, we can significantly enhance the preservation of other non-target concepts, thereby achieving precise unlearning.
@@
  • To the best of our knowledge, we are the first to propose an unlearning-based concept erasing method for T2V models.
  • We propose negatively-guided velocity prediction augmented with prompt augmentation, mask-based localization regularization, and concept preservation regularization, to achieve robust and precise concepts erasing for T2V models.
  • Extensive experiments demonstrate that our method significantly outperforms current concept erasing techniques across diverse prompt types and video generation models.
@@ 2 Related Work
While U-Net-based approaches achieved early success, they struggle with modeling long-range temporal dependencies and suffer from computational inefficiencies. To address these limitations, recent T2V models have shifted toward Diffusion Transformer (DiT) [24] architectures and removed cross-attention. CogVideoX [31] adopts a Multimodal DiT (MMDiT)  [10] that concatenates text and visual tokens as input to a cross-modal 3D full-attention layer, improving both efficiency and performance. HunyuanVideo [21] further explores this direction by introducing dual-stream and single-stream DiT variants: the former processes text and visual tokens separately, while the latter feeds their concatenation into a full-attention layer.
@@
The second category is unlearning, which alters the model weights by zero-shot model editing or fine-tuning methods. Zero-shot approaches directly manipulate the model weights—typically within the cross-attention layers—to suppress undesired concepts, as in UCE [13] and RECE [14]. While efficient and easy to apply, these methods rely heavily on cross-attention mechanism, making them incompatible with recent T2V architectures. In contrast, fine-tuning-based methods offer a more generalizable solution. ESD [12] erases concepts by learning to predict noise guided by negative prompts. AdvUnlearn [33] and Receler [18] build upon ESD with adversarial strategies to improve erasure robustness. Despite the progress in the T2I domain, experiments (see Section 4.1) show that existing methods fail to achieve robust and precise concept erasure in the T2V setting.
@@ 3 Method
Figure 2: Overview of T2VUnlearning pipeline. T2VUnlearning performs unlearning by inserting adapters after each cross-modal full-attention layer. It generates pseudo data through prompt augmentation, and fine-tunes the model using an unlearning loss $\mathcal{L}_{unlearn}$ to erase the target concept, a mask-based localization regularization $\mathcal{L}_{loc}$, and a preservation regularization $\mathcal{L}_{pre}$ to retain non-target concepts.
@@ 3.1 Background
Diffusion models [16] learn to generate data by reversing a process that gradually adds Gaussian noise to input. Given an input data point $\mathbf{x}$ and condition $y$, a diffusion model $\epsilon_{\theta}$ is trained to predict the noise $\epsilon$ from a noisy sample $\mathbf{x}_{t}$ at a given time step $t$:
@@
In T2V diffusion models, the training objectives and generation paradigms often differ. In this work, we focus on two widely used open-source models: CogVideoX [31] and HunyuanVideo [21].
CogVideoX employs v-prediction parameterization [26], where the model is trained to predict a velocity vector. Specifically, given a video latent representation $\mathbf{x}$ and condition $y$, the model adds Gaussian noise $\epsilon$ by $t$ steps and the noisy sample can be expressed as $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon$, where $\alpha_{t}$ and $\sigma_{t}$ are scheduling coefficients. The model $\mathbf{v}_{\theta}$ is encouraged to predict the ground truth velocity vector $\mathbf{v}=\alpha_{t}\mathbf{x}-\sigma_{t}\epsilon$ by the following loss function:
@@
HunyuanVideo is based on rectified flow [23], which trains the model to learn a vector field that maps a complex data distribution to a simple Gaussian distribution. Specifically, given a latent video representation $\mathbf{x}$ and condition $y$, it initializes Gaussian noise $\epsilon$ and constructs a training sample $\mathbf{x}_{t}$ at step $t$ by linear interpolation between $\mathbf{x}$ and $\epsilon$. The model $\mathbf{v}_{\theta}$ is trained to predict the velocity $\mathbf{u}_{t}=d\mathbf{x}_{t}/dt$ that guides the noise back to the data by the following loss function:
$\mathcal{L}_{\textrm{hunyuan}}=\mathbb{E}_{\mathbf{x},y,\mathbf{x}_{0},t}\left\|\mathbf{u}_{t}-\mathbf{v}_{\theta}\left(\mathbf{x}_{t},y,t\right)\right\|_{2}^{2},$ (3)
@@
$P_{\theta^{{}^{\prime}}}(\mathbf{x})\propto\frac{P_{\theta}(\mathbf{x})}{P_{\theta}(c|\mathbf{x})^{\eta}},$ (4)
where the $\theta^{{}^{\prime}}$ indicate new model weights after unlearning, and $\eta$ is a tunable parameter indicating unlearning strength. Increasing $\eta$ forces the model to lower the probability of generating contents containing target concepts, and the target concepts are more likely to be completely erased.
@@ 3.2 Overview of Pipeline
T2VUnlearning is an adapter-based fine-tuning method that achieves robust and precise unlearning through three strategies: negatively-guided velocity prediction with prompt augmentation, mask-based localization regularization, and concept preservation regularization. An overview of the pipeline is illustrated in Figure 2.
3.3 Negatively-guided Velocity Prediction with Prompt Augmentation
Given any T2V model $\mathbf{v}_{\theta}$ with weights $\theta$, our goal is to obtain an unlearned model weight $\theta^{{}^{\prime}}$ that reduces the probability of generating video $\mathbf{x}$ described by target concept $c$. Based on Equation 4, we reformulate the expression in terms of probability density. At diffusion timestep $t$, this leads to:
Using Bayes’ theorem $p(c|{\mathbf{x}_{t}})=\frac{p({\mathbf{x}_{t}}|c)p(c)}{p({\mathbf{x}_{t}})}$, we take the logarithm of both sides and compute the gradient with respect to $\mathbf{x}_{t}$}, rewriting it in the form of score prediction:
$-\eta(\nabla_{\mathbf{x}_{t}}\log p_{\theta}({\mathbf{x}_{t}}|c)-\nabla_{\mathbf{x}_{t}}\log p_{\theta}({\mathbf{x}_{t}})),$
which can be reparameterized to the form of velocity prediction. Specifically, for HunyuanVideo with gaussian probability path $p(\mathbf{x}_{t}|\mathbf{x})=N(\mathbf{x}_{t};(1-t)\mathbf{x},t^{2}I)$, there exists an equivalence [22] between velocity $\mathbf{u}_{t}$ and score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$:
$\mathbf{u}_{t}=-\frac{t}{1-t}\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})-\frac{1}{1-t}\mathbf{x}_{t}.$ (7)
This formula implies that knowing $\mathbf{u}_{t}$ allows us to compute $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$, and consequently, the score estimator $\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t})$ can be equivalently converted to the velocity estimator $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$. Similarly, for CogVideoX, given the diffusion noise assumptions and Tweedie’s formula [9], we can show that the score estimator $\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t})$ can be converted to the velocity estimator $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ (see Appendix 2). Therefore, the objective in Equation 6 can be reparameterized to velocity prediction:
@@ 3.3 Negatively-guided Velocity Prediction with Prompt Augmentation
$\mathcal{L}_{unlearn}$ $=\mathbb{E}_{\mathbf{x},c,\epsilon,t}\left\|\mathbf{v}_{neg}-\mathbf{v}_{\theta^{{}^{\prime}}}\left(\mathbf{x}_{t},c,t\right)\right\|_{2}^{2},$
where $\mathbf{v}_{neg}$ is the right-hand side of Equation 6 rewritten in terms of velocity. Note that $\mathbf{v}_{neg}$ is similar to negative guidance in classifier-free guidance, thus we refer to it as negatively-guided velocity prediction.
Inspired by ESD [12], we use the original T2V model to generate pseudo data instead of manually collecting real training data. Specifically, $\mathbf{x}_{t}$ at step $t$ is obtained by denoising a sampled Gaussian noise to timestep $t$ instead of adding noise to real data. Previous works generate $\mathbf{x}_{t}$ using only the target concept as the prompt. However, training solely on simple instances of concept $c$ is insufficient to reduce the generation probability for contents with additional context or detail, because these more complex examples come from a different distribution that the model hasn’t seen during fine-tuning.
Figure 3: Example of prompt augmentation.
To address this problem, we propose a simple yet effective solution: augmenting the prompts used to generate $\mathbf{x}_{t}$. As shown in Figure 3, we followed the prompt-rewriting strategies employed during the training of CogVideoX, utilizing an uncensored variant of LLaMA3***https://huggingface.co/Orenguteng/Llama-3-8B-Lexi-Uncensored and apply in-context prompting to generate augmented prompts that align with the format requirements of T2V models. Experimental results (see Section 4.4) demonstrate that this approach achieves robust unlearning when tested against a wide range of paraphrased prompts, including those written by humans.
@@ 3.4 Mask-based Localization
Incorporating contextual information into the pseudo data $\mathbf{x}$ can inadvertently cause the model to forget knowledge of those additional contexts and details. Prior work in T2I, such as  Huang et al. [18], addresses this issue by computing a concept-related cross-attention mask and restricting model updates to the masked region. Despite the absence of explicit cross-attention in T2V models, interactions between text and visual tokens still exist. As demonstrated by DiTCtrl [3], the query-key (QK) maps in full-attention layers can be divided into distinct regions, among which the text-to-video region exhibits interactions analogous to cross-attention. Consequently, extracting a concept-related attention mask based on the text-to-video region of the QK map is feasible (as shown in Figure 4).
Figure 4: Example of obtaining attention mask through full-attention QK map.